OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (whpalmer4):

point-slope form is \[y-y_0 = m(x-x_0)\] where \(m\) is slope and \((x_0,y_0)\) is a point on the line. We can find \(m\) from the formula for slope given two points \((x_1,y_1),(x_2,y_2)\) which is \[m = \frac{y_2-y_1}{x_2-x_1}\] Plug and chug! :-)

OpenStudy (anonymous):

flutter boi